### v1

# thanks: Corresponding author: Sebastian Schulz

Abstract
We present Learning-Driven Annealing (LDA), a framework that links individual quantum annealing evolutions into a global solution strategy to mitigate hardware constraints such as short annealing times and integrated control errors. Unlike other iterative methods, LDA does not tune the annealing procedure (e.g. annealing time or annealing schedule), but instead learns about the problem structure to adaptively modify the problem Hamiltonian. By deforming the instantaneous energy spectrum, LDA suppresses transitions into high-energy states and focuses the evolution into low-energy regions of the Hilbert space. We demonstrate the efficacy of LDA by developing a hybrid quantum-classical solver for large-scale spin glasses. The hybrid solver is based on a comprehensive study of the internal structure of spin glasses, outperforming other quantum and classical algorithms (e.g., reverse annealing, cyclic annealing, simulated annealing, Gurobi, Toshiba’s SBM, VeloxQ and D-Wave hybrid) on $5580$-qubit problem instances in both runtime and lowest energy. LDA is a step towards practical quantum computation that enables today’s quantum devices to compete with classical solvers.
Quantum computing, Quantum annealing, Learning-driven annealing, Feature Hamiltonian, Hybrid quantum-classical optimization, Spin glass
I Introduction
Combinatorial optimization problems (COPs) are ubiquitous in computer science, with important applications in finance [COP_Finance], scheduling [COP_Scheduling], machine learning [COP_ML_1, COP_ML_2], computational biology [COP_Biology], and operations research [COP_OperationsResearch]. For many such problems, finding the optimal solution is equivalent to finding the ground state of an associated Ising spin-glass system [ISING_Formulations]. The hardness of COPs is related to the presence of opposing spin interactions, leading to frustration and a glass phase. The latter is characterized by the presence of many low-energy local minima, that are separated by large energy barriers. This makes solving large-scale COPs often intractable on classical computers, quickly exceeding runtimes of $24$ hours.
With the advent of new generations of quantum annealers from D-Wave Systems Inc. comprising more than $5000$ qubits, a promising approach is the use of quantum annealing (QA) [QA_Industry, QA_Supremacy, ThreeDWaveGenerations]. Inspired by the cooling of physical systems, QA uses quantum fluctuations caused by a transverse field to navigate the energy landscape of the spin system [QA_Overview]. If the quantum evolution is performed adiabatically, the system arrives in the ground state of the spin glass. However, if the evolution is too fast, the system scatters into higher energy states through a series of quantum phase transitions (QPTs, used in the sense of Refs. [QA_QPT_1, QA_QPT_2, QA_QPT_3, SG_QPT_Theory]). Hence, on real devices, finite annealing times and integrated control errors limit the performance, often making deep low-energy states (including the ground state) inaccessible [QA_fails_1, QA_fails_2]. To address this issue, several authors recently investigated iterative procedures, such as reverse annealing [IRA_Theory_1, IRA_Theory_2, IRA_Theory_3, IRA_Thermal_1] and cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2, Ra_Hg], in which the annealing is repeatedly initialized from the best known classical state. In doing so, the hope is that QA gradually converges to the ground state, but in practice, these procedures often fail, as they become stuck in high-energy valleys [IRA_Failing_1, IRA_Failing_2].
We show that modifications to the problem Hamiltonian itself can greatly improve the performance of iterative QA on noisy hardware. With Learning-Driven Annealing (LDA), we propose a framework to link individual QA runs into a global solution strategy to mitigate hardware constraints such as finite annealing times and integrated control errors [QPU_ICE]. LDA works by analyzing the states sampled from QA to adaptively modify the problem Hamiltonian to the information learned about the energy landscape. This energetically isolates low-energy valleys in the instantaneous energy spectrum, focussing subsequent annealing runs into a promising region of the Hilbert space. As a result, the QPU can reach deeper regions of the energy spectrum that are otherwise inaccessible, enabling NISQ devices to tackle any COP that can be embedded onto their hardware.
We demonstrate the capabilities of LDA by designing a hybrid quantum-classical solver for large-scale spin glasses, that is based on alternating local and global search protocols. Using the D-Wave Advantage 5.4 QPU [QPU_Advantage_5_4] in Jülich, Germany, we benchmark our hybrid solver against leading quantum and classical algorithms, including reverse annealing [IRA_Theory_1], cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2], simulated annealing using JUPTSA [JUPTSA], Gurobi [Gurobi], Toshiba’s SBM [Toshiba_SBM], VeloxQ [VeloxQ] and D-Wave hybrid [D_Wave_Hybrid], on $5580$-qubit NAT-7 [NAT_7_1, NAT_7_2] spin glasses. The results show that our hybrid solver outperforms all competing strategies in both runtime and lowest energy. LDA is a step towards practical quantum computation that enables today’s quantum annealers to compete with classical solvers.
The remainder of this paper is organized as follows: In Sec. II, we introduce the spin-glass formulation and provide insights into its underlying structures. Section LABEL:sec:qa concerns adiabatic quantum evolutions and discusses its limitations on modern quantum annealers. In Sec. LABEL:sec:lda, we introduce the LDA framework, showcasing how learned information from sampled states can be used to guide the annealing evolution by modifying the problem Hamiltonian. In Sec. LABEL:sec:hybrid_optimization, we demonstrate how LDA can be used in advanced algorithms by designing a hybrid solver for spin glasses based on alternating local and global search protocols. Finally, in Sec. LABEL:sec:benchmarks, we benchmark our hybrid solver against various classical and quantum solvers, followed by a summary of our findings in Sec. LABEL:sec:conclusion.
Figure 1: A $24$-qubit spin-glass with $N_{dom}=3$. (a) Spin-glass graph with the biases $h_{i}$ as nodes and couplers $J_{ij}$ as edges. Dashed edges indicate periodic boundary conditions. The colored rectangles represent the spin domains: $\mathcal{D}_{1}$ (green), $\mathcal{D}_{2}$ (purple), and $\mathcal{D}_{3}$ (orange). Each domain has two distinct minima: $\alpha$ and $\overline{\alpha}$ ($\mathcal{D}_{1}$), $\beta$ and $\overline{\beta}$ ($\mathcal{D}_{2}$), $\gamma$ and $\overline{\gamma}$ ($\mathcal{D}_{3}$). Panels (a.1-3) show the eigenenergies $E$ of the corresponding domains as a function of the Hamming distance $d_{ham}$ to the respective ground state $\alpha,\beta,\gamma$. (b) Hypercube representation of the Hilbert space, with each axis representing the Hamming distance $d_{ham}$ from the ground state ($\alpha,\beta,\gamma$) of the respective spin domains $\mathcal{D}_{k}$. Color indicates the lowest eigenenergy at each point. The corners of the hypercube correspond to the $82^{N_{dom}}=$ 8 local minima of the spin glass. (c) Energy landscape of the spin glass obtained by unfolding the hypercube facets. Each point is projected onto the closest facet (smallest Hamming distance), with the blue surface representing the lowest eigenenergies on these facets. The eight local minima are marked by colored diamonds at the facets’ corners, with the states surrounding it being the energy valleys. Additionally, black dots indicate $2000$ samples from a quantum annealing simulation with annealing time $T=10$, with the probability density shown on the lower orange plane. Panels (c.1-3) highlight the specific biases $h_{i}$ and couplers $J_{ij}$ satisfied by three of the eight local minima, with the unsatisfied terms set to 00 (white). The color scheme matches that of panel (a).
II Spin-glass problems
Spin glasses (SG) are magnetic systems of quenched disorder, imposing conflicts between interacting magnetic moments [SG_Theory]. In the Edwards-Anderson model of SG [SG_Theory_EA], $N$ Ising spins $\sigma^{z}_{i}$, $i\in\{0,\,\dots\,,\,N-1\}$, are placed on the sites of a regular graph, with the Hamiltonian defined as
$H=\sum_{i<j}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{z}_{i}.$ (1)
Here $J_{ij}$ with $i<j$ denotes the coupling strength between spins $i$ and $j$, and $h_{i}$ represents a locally applied external magnetic field bias on spin $i$.
We consider both triangular lattices with periodic boundary conditions (see Fig. 1a) and the Pegasus graph used by the D-Wave Advantage 5.4 quantum annealer [QPU_Advantage_5_4]. In the following, we choose |$\lvert$|

### v2

@@
 LDA is a step towards practical quantum computation that enables today’s quantum devices to compete with classical solvers.
+[…]
 Quantum computing, Quantum annealing, Learning-driven annealing, Feature Hamiltonian, Hybrid quantum-classical optimization, Spin glass
-I Introduction
+1 Introduction
 Combinatorial optimization problems (COPs) are ubiquitous in computer science, with important applications in finance [COP_Finance], scheduling [COP_Scheduling], machine learning [COP_ML_1, COP_ML_2], computational biology [COP_Biology], and operations research [COP_OperationsResearch].
@@
 In doing so, the hope is that QA gradually converges to the ground state, but in practice, these procedures often fail, as they become stuck in high-energy valleys [IRA_Failing_1, IRA_Failing_2].
-We show that modifications to the problem Hamiltonian itself can greatly improve the performance of iterative QA on noisy hardware.
+We show that systematically learned modifications to the problem Hamiltonian itself can greatly improve the performance of iterative QA on noisy hardware.
 With Learning-Driven Annealing (LDA), we propose a framework to link individual QA runs into a global solution strategy to mitigate hardware constraints such as finite annealing times and integrated control errors [QPU_ICE].
@@
 This energetically isolates low-energy valleys in the instantaneous energy spectrum, focussing subsequent annealing runs into a promising region of the Hilbert space.
-As a result, the QPU can reach deeper regions of the energy spectrum that are otherwise inaccessible, enabling NISQ devices to tackle any COP that can be embedded onto their hardware.
+As a result, the QPU can reach deeper regions of the energy spectrum that are otherwise inaccessible.
+While we do not claim a quantum advantage per se, LDA enables current NISQ devices to tackle COPs more efficiently as long as they can be embedded onto their hardware.
 We demonstrate the capabilities of LDA by designing a hybrid quantum-classical solver for large-scale spin glasses, that is based on alternating local and global search protocols.
-Using the D-Wave Advantage 5.4 QPU [QPU_Advantage_5_4] in Jülich, Germany, we benchmark our hybrid solver against leading quantum and classical algorithms, including reverse annealing [IRA_Theory_1], cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2], simulated annealing using JUPTSA [JUPTSA], Gurobi [Gurobi], Toshiba’s SBM [Toshiba_SBM], VeloxQ [VeloxQ] and D-Wave hybrid [D_Wave_Hybrid], on $5580$-qubit NAT-7 [NAT_7_1, NAT_7_2] spin glasses.
+Using the D-Wave Advantage1 5.4 QPU [QPU_Advantage_5_4] in Jülich, Germany, we benchmark our hybrid solver against leading quantum and classical algorithms, including reverse annealing [IRA_Theory_1], cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2], simulated annealing using JUPTSA [JUPTSA], Gurobi [Gurobi], Toshiba’s SBM [Toshiba_SBM], VeloxQ [VeloxQ, VeloxQ2] and D-Wave hybrid [D_Wave_Hybrid], on $5580$-qubit NAT-7 [NAT_7_1, NAT_7_2] spin glasses.
 The results show that our hybrid solver outperforms all competing strategies in both runtime and lowest energy.
@@
 The remainder of this paper is organized as follows: In Sec.
-II, we introduce the spin-glass formulation and provide insights into its underlying structures.
+2, we introduce the spin-glass formulation and provide insights into its underlying structures.
 Section LABEL:sec:qa concerns adiabatic quantum evolutions and discusses its limitations on modern quantum annealers.
@@
 Color indicates the lowest eigenenergy at each point.
-The corners of the hypercube correspond to the $82^{N_{dom}}=$ 8 local minima of the spin glass.
+The corners of the hypercube correspond to the $2^{N_{dom}}=8$ local minima of the spin glass.
 (c) Energy landscape of the spin glass obtained by unfolding the hypercube facets.
@@
 Additionally, black dots indicate $2000$ samples from a quantum annealing simulation with annealing time $T=10$, with the probability density shown on the lower orange plane.
-Panels (c.1-3) highlight the specific biases $h_{i}$ and couplers $J_{ij}$ satisfied by three of the eight local minima, with the unsatisfied terms set to 00 (white).
+Panels (c.1-3) highlight the specific biases $h_{i}$ and couplers $J_{ij}$ satisfied by three of the eight local minima, with the unsatisfied terms set to 0 (white).
 The color scheme matches that of panel (a).
-II Spin-glass problems
+2 Spin-glass problems
 Spin glasses (SG) are magnetic systems of quenched disorder, imposing conflicts between interacting magnetic moments [SG_Theory].
-In the Edwards-Anderson model of SG [SG_Theory_EA], $N$ Ising spins $\sigma^{z}_{i}$, $i\in\{0,\,\dots\,,\,N-1\}$, are placed on the sites of a regular graph, with the Hamiltonian defined as
+In the Edwards-Anderson model of SG [SG_Theory_EA], NN Ising spins $\sigma^{z}_{i}$, $i\in\{0,\,\dots\,,\,N-1\}$, are placed on the sites of a regular graph, with the Hamiltonian defined as
 $H=\sum_{i<j}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{z}_{i}.$ (1)
-Here $J_{ij}$ with $i<j$ denotes the coupling strength between spins $i$ and $j$, and $h_{i}$ represents a locally applied external magnetic field bias on spin $i$.
+Here $J_{ij}$ with $i<j$ denotes the coupling strength between spins $i$ and $j$, and $h_{i}$ represents a locally applied external magnetic field bias on spin ii.
 We consider both triangular lattices with periodic boundary conditions (see Fig.
-1a) and the Pegasus graph used by the D-Wave Advantage 5.4 quantum annealer [QPU_Advantage_5_4].
-In the following, we choose |$\lvert$|
+1a) and the Pegasus graph used by the D-Wave Advantage1 5.4 quantum annealer [QPU_Advantage_5_4].
+In the following, we choose $\lvert$